# 11footnotetext: Equal Contribution22footnotetext: Equal Advising Contribution

### v1

@@ 1 Introduction
However, the best performing modern LLMs are usually massive in terms of the parameter counts (e.g., 175B parameters for GPT3 (Brown et al., 2020) and 340B for PaLM2 (Anil et al., 2023)), and are therefore difficult to train with limited resources. This motivates a practical question, that to the best of our knowledge has not been addressed systematically, and thus we address it in this article: what is the best embedding model one can train from a backbone decoder-only LLM with a fixed budget?
@@
Contribution
We comprehensively experimented with contrastively fine-tuning language models into text embedding models. We analysed how the choice of model sizes, data quantities, and fine-tuning methods affect the performance in the resource-constrained regime. We compiled the findings into an algorithm that, given a fixed computational budget, predicts the optimal network architecture, data quantity, and parameter-efficient fine-tuning hyperparameters. We open-source the code to train and evaluate our models at: https://github.com/SeqDM/Efficient-Embeddings.
@@ 2 Related work
Embedding models
Training neural networks to represent text as continuous vectors was popularised by word2vec (Mikolov et al., 2013), which produced semantically meaningful embeddings of words.
@@
Embedding models are versatile in their applications. Therefore, a broad and diverse benchmark to provide a robust measure of their performance is called for. The first such benchmark was BEIR introduced by Thakur et al. (2021). It has nine different information retrieval tasks (e.g., duplicate-question retrieval or citation-prediction) on 18 datasets. Recently, Muennighoff et al. (2023b) introduced MTEB (Massive Text Embedding Benchmark), which is substantially larger than BEIR and, in addition to information retrieval, has seven other types of tasks (e.g., clustering, summarisation, and pair classification) using 58 datasets and covering 112 languages. The authors also provide a leaderboard that currently contains more than 220 entries.
Scaling laws
@@
Scaling laws predict the performance of models at a larger scale (e.g., more parameters, more data) from experiments at smaller scales. Kaplan et al. (2020) and Hoffmann et al. (2022) used empirically fitted laws to predict the performance of neural language models at different model sizes and data scales, making it possible to calculate the optimal configuration before training. Muennighoff et al. (2023a) inspected model training in data-constrained, multi-epoch regime. Frantar et al. (2023) looked at scaling laws for language models in the context of weight sparsity. Zhang et al. (2024) investigates the scaling laws for model fine-tuning, including parameter efficient methods: LoRA (Hu et al., 2022) and prompt tuning (Lester et al., 2021). They show that the scaling laws and the best fine-tuning method is task dependent. Biderman et al. (2024) investigates the properties of fine-tuning models with LoRA in details, showing that LoRA underperforms full fine-tuning when fine-tuning the models on mathematical or code datasets.
@@
However, none of these works focus on scaling laws for training embedding models. Our scaling law is targeted at the process of contrastive fine-tuning for achieving a model yielding good embeddings. Moreover, our main goal is to find which strategy for achieving good embeddings is optimal in a budget-restricted settings, while taking into account the popularity of applying parameter efficient methods for fine-tuning, like LoRA or partial model freezing. In spirit, our work is similar to AutoML (He et al., 2021b), which aims to automatically find the optimal ML architecture. However, our method goes beyond and targets the data and the training method as well.
Parameter-efficient fine-tuning
@@
Modern language models have a lot of parameters, and fine-tuning them using consumer-grade hardware is difficult in general. Notwithstanding the recent turn to in-context learning, fine-tuning is still necessary for most applications. A range of techniques have been developed recently, to make the fine-tuning more efficient (Houlsby et al., 2019; Li and Liang, 2021; Qi et al., 2022; He et al., 2021a). These approaches are all applicable to embedding models. Muennighoff (2022) explored a simple parameter-efficient approach to repurpose GPT models into embedding models where only the bias tensors of the transformer model are updated (Zaken et al., 2022). However, a systematic study of what parameter-efficient methods are optimal under what scenarios has not been performed, which is the aim of our work.
@@ 3.1 Scaling laws and compute-optimal models
The constrained optimisation problem that we aim to solve is the following: given a fine-tuning corpus and a family of pre-trained decoder-only language models of different sizes, minimise the contrastive loss subject to a fixed computational budget.
@@ 3.2 Extracting representations from transformers
Some works apply the last token pooling approach (Neelakantan et al., 2022) instead of using mean pooling described above. Here, however, we default to mean pooling since it is a recently more popular method overall (Wang et al., 2022; Li et al., 2023; Günther et al., 2023).
@@ 3.3 Contrastive fine-tuning
Hard negatives in contrastive learning
In our contrastive learning setting, we use in-batch examples as negatives. Some datasets additionally include tailored negative examples (hard negatives) for each positive datapoint, which can be incorporated into the contrastive loss to promote learning more precise representations.
However, works that do rely on hard negatives often arrive at embeddings by having two distinct training phases: the expensive “unsupervised” phase one involving vast data from the internet, and then a “supervised” phase two, often targeted towards a specific downstream task, where training data of lower quantity – but containing high-quality hard negatives – is used (Li et al., 2023; Wang et al., 2022; Xiao et al., 2023). Since the first phase is more expensive, it will benefit more from scaling laws such as the ones we derive.
@@ 3.4 Fine-tuning methods
Block freezing
In this approach, we freeze the LLM token-embedding layer as well as the first $k$ transformer blocks so that their parameters stay fixed during fine-tuning. Only the latter part of the model is back-propagated through and updates its parameters. To process the same amount of data, this is more computationally efficient than full fine-tuning. By varying the number of frozen blocks, one trades-off between the computational efficiency and flexibility of the model.
Low-rank adaptation (LoRA)
This method was introduced by Hu et al. (2022) to significantly reduce the number of trainable parameters of a neural network while maintaining high performance. LoRA introduces a small number of weights into the model and only trains those. It can be applied to a subset of dense layers of a model. It works by parameterising each dense layer $W$ from a subset of linear layers of a model by two low-rank matrices $A$ and $B$, and using $W+AB$ instead of $W$, while training only $A$ and $B$.
Bias-only tuning
In this method, only bias parameters are fine-tuned, and the rest of the model remains frozen. This approach was proposed by Zaken et al. (2022) as a parameter-efficient fine-tuning method for BERT encoders, and recently applied by Muennighoff (2022) to contrastively fine-tune GPT models.
@@ 4 Experiments
Here, we first specify the relevant details of our experimental setup (Section 4.1). Next, we present the results of our experiments where we contrastively train a grid of models of different sizes, using different computational budgets, and apply different compute-optimal fine-tuning methods with varying hyperparameters (Section 4.2). Based on the collected data, we fit a mathematical formula describing the observed scaling laws (Section 4.3). Finally, we provide general observations and recommendations for efficient contrastive fine-tuning (Section 4.5). [^2] The data collected in our experiments and their interactive visualisations are available in this Colab notebook: https://colab.research.google.com/drive/1EC2QdVrOIXhamLJI51CUywWJ33sApXX8
@@ 4.1 Experimental setup
We use the AdamW optimiser (Loshchilov and Hutter, 2019) and a cosine learning rate scheduler during training. The learning rate first goes through a linear warm-up phase of $1/10$ of the total steps to a peak that is $1/10$ of the pre-training peak learning rate, that is, to learning rates between $1.2\times 10^{-5}$ and $6\times 10^{-5}$. Then, it decays in a cosine schedule to $1/10$ of the maximum at the end of training. We employ gradient checkpointing (Chen et al., 2016) and GradCache (Gao et al., 2021b) to limit the peak GPU RAM usage during training. We use the Transformers (Wolf et al., 2020) library for the training of the models and use the Accelerate (Gugger et al., 2022) library to facilitate multi-GPU training. The batch size in our main line of experiments is 1024 sequences and the context length is 75 tokens. We give the full list of training hyperparameters in Appendix D.
We also perform a less extensive grid of experiments to confirm that our findings are robust to change in hyperparameters. We find that our scaling law formula fitted to losses from models with batch size 1024 and context length of 75 tokens describes the loss for models trained with batch size 2048 and context length of 512 tokens very well. The detailed description and plots are shown in Appendix  F.
@@ 4.2 Experimental results for different methods
Full fine-tuning
With full fine-tuning, the numbers of forward and backward parameters are the same, and the computational cost is straightforwardly $C=6N_{F}D$. In Figure 2(a), we present the final contrastive loss vs. the number of parameters in the network on a log-log scale.
@@
Block freezing
Since the Pythia models we experimented with have $6,12,16,24$ or $32$ blocks, we freeze $\frac{0}{6},\frac{1}{6},\ldots,\frac{5}{6}$ of their blocks if their number is divisible by $6$, and $\frac{0}{8},\frac{1}{8},\ldots,\frac{7}{8}$ if it is divisible by $8$. In this setup, unlike for full fine-tuning, the token-embedding parameters always remain frozen even when all the transformer blocks are active. We have $N_{B}=N_{U}<N_{F}$, and hence cost $C=2N_{F}D+4N_{B}D$.
@@
Bias-only tuning
Here we simply update only the bias parameters and not the weights of the model. Figure 4(a) shows the final training loss across all the model sizes and computational budgets. Increasing the computational budget decreases the optimal loss that is achievable, although the absolute values of the losses are high compared to other fine-tuning methods (the lowest achievable loss is above 0.4 with bias-only tuning).
@@
Low-rank adaptation
We use the LoRA implementation from the PEFT library (Mangrulkar et al., 2022). We attach the adapter to all the dense layers of the network, and vary the adapter rank from $8$ up to $2048$, since even lower ranks resulted in an unstable training.
@@
Lookup table of optimal model sizes for different methods
In Figure 6(a) (for full fine-tuning), Figure 6(b) (for LoRA), and Figure 12 (for freezing and bias tuning, in Appendix A), we visualise the relationship between the optimal model size and the computational budget. This makes looking up the optimal configuration for a given method at a given budget very convenient.
Downstream performance
In Figures 7, 8, 9, and 10, we present the IsoFLOP plots of the downstream performance for all the four fine-tuning methods on the subset of the MTEB benchmark (described in Appendix B). Based on the performance results, we also present the optimal size choices per budget in Figure 13 (which is a performance analogue of Figure 12) and the optimal performance plot in Figure 11 (a performance analogue of Figure 1). These are mostly consistent with the loss results, albeit more noisy. In Appendix C, we demonstrate the significant correlation between the final training loss and the downstream task performance.
@@ 4.3 Scaling laws for embeddings
where $S$ is an additional variable representing the trainable fraction of parameters, and $a_{d},b_{d},a_{s},b_{s},c_{s}\in\mathbb{R}$ are additional coefficients to be fitted. We check that this new formula fits the data better than the original one, and conjecture that it can be used for larger models and larger computational budgets. More details about the derivation are presented in Appendix E.
@@ 4.4 Compute-optimal frontier and recipe
Figure 6: Optimal model size vs. computational budget for full fine-tuning (a), and LoRA (b).
@@ 4.5 Takeaways
We compile the takeaways from the extensive experiments, by analysing each fine-tuning method individually and contrasting them afterwards.
@@ 5 Limitations and future work
Other families of models
In this paper, we experimented with the Pythia (Biderman et al., 2023) suite of pre-trained models, which are of different sizes but have all been trained on 300B tokens. This means that they have been over-trained with respect to the Chinchilla scaling law (Hoffmann et al., 2022) to different degrees. The relationship between the optimal model size and the computational budget might differ for a suite of pre-trained models with different over-training ratios. However, we notice that prior efforts (Dettmers et al., 2024) have found that fine-tuning conclusions reached with Pythia do generalise to other model families such as OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023), and BLOOM (Scao et al., 2022), in the quantised setting.
Averaging
Due to constraints on our computational resources, we ran each experiment exactly once. Averaging over multiple random seeds for the experiments would reduce the variance involved in each, and potentially result in less noisy conclusions. This is a straightforward extension that we leave to carry out in future work, should we have more resources. We also plan to evaluate the trained models on the entirety of the MTEB benchmark (Muennighoff et al., 2023b), instead of using the losses from the last steps of training, to better reflect the performances of embedding models.
Other forms of embedding readout
We only experimented with mean pooling as a way of extracting embeddings from the transformer models. Other readout methods could be used, like the max pooling or extracting the hidden state corresponding to the end-of-sequence token, which might result in different scaling laws and optimal frontiers.

### v2

# Repurposing Language Models into Embedding Models: Finding the Compute-Optimal Recipe

@@ 1 Introduction
However, the best performing modern LLMs are usually massive in terms of the parameter counts (e.g., 175B parameters for GPT3 (Brown et al., 2020) and 340B for PaLM2 (Anil et al., 2023)), and are therefore difficult to train with limited resources. This motivates a practical question, that to the best of our knowledge has not been addressed systematically, and thus we address it in this paper: what is the best embedding model one can train from a backbone decoder-only LLM with a fixed training compute budget?
@@
Contribution   We comprehensively experimented with contrastively fine-tuning language models into text embedding models. We analysed how the choice of model sizes, data quantities, and fine-tuning methods affect the performance in the resource-constrained training regime. We compiled the findings into an algorithm that, given a fixed computational budget, predicts the optimal network architecture, data quantity, and parameter-efficient fine-tuning hyperparameters. We open-source the code to train and evaluate our models at: https://github.com/SeqDM/Efficient-Embeddings.
@@ 2 Related work
Embedding models   Training neural networks to represent text as continuous vectors was popularised by word2vec (Mikolov et al., 2013), which produced semantically meaningful embeddings of words.
@@
Benchmarks for embedding models   Embedding models are versatile in their applications. Therefore, a broad and diverse benchmark to provide a robust measure of their performance is called for. The first such benchmark was BEIR introduced by Thakur et al. (2021). It has nine different information retrieval tasks (e.g., duplicate-question retrieval or citation-prediction) on 18 datasets. Recently, Muennighoff et al. (2023b) introduced MTEB (Massive Text Embedding Benchmark), which is substantially larger than BEIR and, in addition to information retrieval, has seven other types of tasks (e.g., clustering, summarisation, and pair classification) using 58 datasets and covering 112 languages. The authors also provide a leaderboard that currently contains more than 220 entries.
@@
Scaling laws   Scaling laws predict the performance of models at larger scale (e.g., more parameters, more data) from experiments at smaller scales. Kaplan et al. (2020) and Hoffmann et al. (2022) used empirically fitted laws to predict the performance of neural language models at different model sizes and data scales, making it possible to calculate the optimal configuration before training. Muennighoff et al. (2023a) inspected model training in data-constrained, multi-epoch regime. Frantar et al. (2023) looked at scaling laws for language models in the context of weight sparsity. Zhang et al. (2024) investigates the scaling laws for model fine-tuning, including parameter efficient methods: LoRA (Hu et al., 2022) and prompt tuning (Lester et al., 2021). They show that the scaling laws and the best fine-tuning method are task dependent. Biderman et al. (2024) investigates the properties of fine-tuning models with LoRA in detail, showing that LoRA underperforms full fine-tuning when fine-tuning the models on mathematical or code datasets.
@@
The single most related investigation is a concurrent work Fang et al. (2024), where the scaling laws for encoder models for retrieval are investigated. There are significant differences in the settings we consider. Our focus is on investigating the process of fine-tuning decoder-only models for good quality embeddings. Moreover, our main goal is to find which strategy for achieving good embeddings is optimal in a budget-restricted settings, while taking into account the popularity of applying parameter efficient methods for fine-tuning, like LoRA or partial model freezing. In spirit, our work is similar to AutoML (He et al., 2021b), which aims to automatically find the optimal ML architecture. However, our method goes beyond and targets the data and the training method as well.
@@
Parameter-efficient fine-tuning   Modern language models have a lot of parameters, and fine-tuning them using consumer-grade hardware is difficult in general. Notwithstanding the recent turn to in-context learning, fine-tuning is still necessary for most applications. A range of techniques have been developed recently, to make fine-tuning more efficient (Houlsby et al., 2019; Li and Liang, 2021; Qi et al., 2022; He et al., 2021a). These approaches are all applicable to embedding models. Muennighoff (2022) explored a simple parameter-efficient approach to repurpose GPT models into embedding models where only the bias tensors of the transformer model are updated (Zaken et al., 2022). (Sun et al., 2024) developed a parameter-efficient method designed specifically for fine-tuning embedding models. However, a systematic study of what parameter-efficient methods are optimal under what scenarios has not been performed, which is the aim of our work.
@@ 3.1 Scaling laws and compute-optimal models
The constraint optimisation problem that we aim to solve is the following: given a fine-tuning corpus and a family of pre-trained decoder-only language models of different sizes, minimise the contrastive loss subject to a fixed computational budget.
@@ 3.2 Extracting representations from transformers
Some works apply the last token pooling approach (Neelakantan et al., 2022) instead of using mean pooling described above. Here, however, we default to mean pooling as in (Wang et al., 2022; Li et al., 2023; Günther et al., 2023). It is a more popular method overall, and moreover, we find it to yield better performance, which we demonstrate in Appendix E.3.
@@ 3.3 Contrastive fine-tuning
Hard negatives in contrastive learning   In our contrastive learning setting, we use in-batch examples as negatives. Some datasets additionally include tailored negative examples (hard negatives) for each positive datapoint, which can be incorporated into the contrastive loss to promote learning more precise representations. Moreover, there are works that focus on approaches for mining hard negatives which result in better training data in the context of specific downstream tasks (Xiong et al., 2021; Zhan et al., 2021).
However, recent works aiming at training powerful, general-purpose embedding models that do rely on datasets with hard negatives often arrive at embeddings by having two distinct training phases: the expensive “unsupervised” phase one involving vast data from the internet, and then a “supervised” phase two, often targeted towards a specific downstream task, where training data of lower quantity – but containing high-quality hard negatives – is used (Li et al., 2023; Wang et al., 2022; Xiao et al., 2023). Since the first phase is more expensive, it will benefit more from scaling laws such as the ones we derive.
@@ 3.4 Fine-tuning methods
Block freezing   In this approach, we freeze the LLM token-embedding layer as well as the first $k$ transformer blocks so that their parameters stay fixed during fine-tuning. Only the latter part of the model is back-propagated through and updates its parameters. To process the same amount of data, this is more computationally efficient than full fine-tuning. By varying the number of frozen blocks, one trades-off between the computational efficiency and flexibility of the model.
Low-rank adaptation (LoRA)   This method was introduced by Hu et al. (2022) to significantly reduce the number of trainable parameters of a neural network while maintaining high performance. LoRA introduces a small number of weights into the model and only trains those. It can be applied to a subset of dense layers of a model. It works by parametrising each dense layer $W$ from a subset of linear layers of a model by two low-rank matrices $A$ and $B$, and using $W+AB$ instead of $W$, while training only $A$ and $B$.
Bias-only tuning   In this method, only bias parameters are fine-tuned, and the rest of the model remains frozen. This approach was proposed by Zaken et al. (2022) as a parameter-efficient fine-tuning method for BERT encoders, and recently applied by Muennighoff (2022) to contrastively fine-tune GPT models.
@@ 4 Experiments
We first specify the relevant details of our experimental setup (Section 4.1). Next, we present the results of our experiments where we contrastively train a grid of models of different sizes, using different computational budgets, and apply different compute-optimal fine-tuning methods with varying hyperparameters (Section 4.2). Based on the collected data, we fit a mathematical formula describing the observed scaling laws (Section 4.3). Finally, we provide general observations and recommendations for efficient contrastive fine-tuning (Section 4.6). In these experiments [^2] The data collected in our experiments and their interactive visualisations are available in this Colab notebook: https://colab.research.google.com/drive/1EC2QdVrOIXhamLJI51CUywWJ33sApXX8 we investigate the compute-constrained setup, but we present limited results on the data-constrained setup in Appendix D.
@@ 4.1 Experimental setup
We use the AdamW optimiser (Loshchilov and Hutter, 2019) and a cosine learning rate scheduler during training. The learning rate first goes through a linear warm-up phase of $1/10$ of the total steps to a peak that is $1/10$ of the pre-training peak learning rate, that is, to learning rates between $1.2\times 10^{-5}$ and $6\times 10^{-5}$. Then, it decays in a cosine schedule to $1/10$ of the maximum at the end of training. We employ gradient checkpointing (Chen et al., 2016) and GradCache (Gao et al., 2021b) to limit the peak GPU RAM usage during training. We use the Transformers (Wolf et al., 2020) library for the training of the models and use the Accelerate (Gugger et al., 2022) library to facilitate multi-GPU training. The batch size in our main line of experiments is 1024 sequences and the context length is 75 tokens. We give the full list of training hyperparameters in Appendix E.
We also perform a less extensive grid of experiments to confirm that our findings are robust to change in hyperparameters. We find that our scaling law formula fitted to losses from models with batch size 1024 and context length of 75 tokens describes the loss for models trained with batch size 2048 and context length of 512 tokens very well. The detailed description and plots are shown in Appendix  G.
@@ 4.2 Experimental results for different methods
Full fine-tuning   With full fine-tuning, the numbers of forward and backward parameters are the same, and the computational cost is straightforwardly $C=6N_{F}D$. In Figure 2(a), we present the final contrastive loss vs. the number of parameters in the network on a log-log scale.
@@
Block freezing   Since the Pythia models we experimented with have $6,12,16,24$ or $32$ blocks, we freeze $\frac{0}{6},\frac{1}{6},\ldots,\frac{5}{6}$ of their blocks if their number is divisible by $6$, and $\frac{0}{8},\frac{1}{8},\ldots,\frac{7}{8}$ if it is divisible by $8$. In this setup, unlike for full fine-tuning, the token-embedding parameters always remain frozen even when all the transformer blocks are active. We have $N_{B}=N_{U}<N_{F}$, and hence cost $C=2N_{F}D+4N_{B}D$.
@@
Bias-only tuning   Here we simply update only the bias parameters and not the weights of the model. Figure 4(a) shows the final training loss across all the model sizes and computational budgets. Increasing the computational budget decreases the optimal loss that is achievable, although the absolute values of the losses are high compared to other fine-tuning methods (the lowest achievable loss is above 0.4 with bias-only tuning).
@@
Low-rank adaptation   We use the LoRA implementation from the PEFT library (Mangrulkar et al., 2022). We attach the adapter to all the dense layers of the network, and vary the adapter rank from $8$ up to $2048$, since even lower ranks resulted in an unstable training.
@@
Lookup table of optimal model sizes for different methods   In Figure 6(a) (for full fine-tuning), Figure 6(b) (for LoRA), and Figure 13 (for freezing and bias tuning, in Appendix A), we visualise the relationship between the optimal model size and the computational budget. This makes looking up the optimal configuration for a given method at a given budget very convenient.
Figure 6: Optimal model size vs. computational budget for (a) full fine-tuning, and (b) LoRA.
Downstream performance   In Figures 8, 9, 10, and 11 (in Appendix A) we present the IsoFLOP plots of the downstream performance for all the four fine-tuning methods on the subset of the MTEB benchmark (described in Appendix B). Based on the performance results, we also present the optimal size choices per budget in Figure 15 in Appendix A (which is a performance analogue of Figure 13), and the optimal performance plot in Figure 12 (a performance analogue of Figure 1). These are mostly consistent with the loss results, albeit more noisy. In Appendix C, we demonstrate the significant correlation between the final training loss and the downstream task performance.
@@ 4.3 Scaling laws for embeddings
where $S$ is an additional variable representing the trainable fraction of parameters, and $a_{d},b_{d},a_{s},b_{s},c_{s}\in\mathbb{R}$ are additional coefficients to be fitted. We check that this new formula fits the data better than the original one, and conjecture that it can be used for larger models and larger computational budgets. More details about the derivation are presented in Appendix F.
@@ 4.4 Compute-optimal frontier and recipe
We briefly discuss the case of data-optimal frontier in Appendix D. We do not emphasise this setup due to it being less standard in language modelling research.
@@
4.5 Generalisation
[…]
Figure 7 presents results for Gemma 2B with both full fine-tuning and LoRA fine-tuning, while results for Gemma 2 2B are shown in Appendix A. In both cases, the formula provides a close approximation of the loss behavior. Although this study is limited to a single model size, it offers an initial indication that our findings may generalize to new models.
Figure 7: Actual loss vs. predicted loss for full fine-tuning and LoRA finetuning. Red points are the models from runs on Pythia suite of size smaller than 2.8B, blue points are from runs on Pythia 2.8B and green points are from runs on Gemma 2B. The formula is fitted using the red points.
@@ 4.6 Takeaways
We compile the takeaways from our extensive experiments by analysing each fine-tuning method individually and contrasting them afterwards.
@@ 5 Limitations and future work
Other families of models   In this paper, we experimented with the Pythia (Biderman et al., 2023) suite of pre-trained models, which are of different sizes but have all been trained on 300B tokens. This means that they have been over-trained with respect to the Chinchilla scaling law (Hoffmann et al., 2022) to different degrees. The relationship between the optimal model size and the computational budget might differ for a suite of pre-trained models with different over-training ratios. We note that prior efforts (Dettmers et al., 2024) have found that fine-tuning conclusions reached with Pythia do generalise to other model families such as OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023), and BLOOM (Scao et al., 2022), in the quantised setting. Indeed, in Section 4.5, we find that a model from a different family (Gemma (Mesnard et al., 2024)) closely fits our loss predictions.
Averaging   Due to constraints on our computational resources, we ran each experiment exactly once. Averaging over multiple random seeds for the experiments would reduce the variance involved in each, and potentially result in less noisy conclusions. This is a straightforward extension that we leave to carry out in future work, should we have more resources. We also plan to evaluate the trained models on the entirety of the MTEB benchmark (Muennighoff et al., 2023b), instead of using the losses from the last steps of training, to better reflect the performances of embedding models.
Other forms of embedding readout   We only experimented with mean pooling as a way of extracting embeddings from the transformer models. Other readout methods could be used, like the max pooling or extracting the hidden state corresponding to the end-of-sequence token, which might result in different scaling laws and optimal frontiers.
Inference cost analysis   In this work, we only focus on the training cost and refrain from taking inference costs into account. However, in certain practical scenarios, the latter may also be relevant.